OpenStudy (lukeamendola):

A = 4πr2. can you help me solve for r

OpenStudy (will.h):

Welcome To OpenStudy

OpenStudy (will.h):

can you show your work so far? if you haven't then how can you solve this

OpenStudy (starlord1):

https://www.mathpapa.com/algebra-calculator.html

OpenStudy (will.h):

StarLord that is totally in accurate he/she wants to solve for r

OpenStudy (starlord1):

oh.... then a=(4)(3.141593)r2 Step 1: Flip the equation. 12.566371r2=a Step 2: Divide both sides by 12.566371. 12.566371r2 12.566371 = a 12.566371 r2=0.079577a Step 3: Take square root. r=√0.079577a or r=−√0.079577a Answer: r=√0.079577a or r=−√0.079577a

563blackghost (563blackghost):

We have no idea as to what `A` or `r` is so we would change this into an equation for r....we would keep `pi` as `the sign for it`.... \(\huge{A= 4 \pi r ^{2}}\) First we would divde by 4.... \(\huge{\frac{A}{4}=\frac{4}{4} \pi r ^{2} \rightarrow \frac{A}{4}= \pi r^{2}}\) Next we would divde \(\Large{\pi}\).... \(\huge{\frac{A}{4} \times \frac{\pi}{\pi} = \frac{\pi}{\pi}r^{2} \rightarrow \frac{A \pi}{4 \pi}=r^{2}}\) Next we would do square root.... \(\huge\frac{\sqrt{A \pi}}{\sqrt{4 \pi}}=\sqrt{r^{2}}\rightarrow r = \frac{\sqrt{A \pi}}{2 \pi}\) That would be your answer :)